Problem 09.020 Benefit/Cost Analysis of a Single Project

A consultant, after 3 months of work, reported that the modified B/C ratio for a city-owned hospital heliport project is 2. If the initial cost is $1.9 million and the annual benefits are $115,000, what is the amount of the annual M&O costs used in the calculation? The report stated that a discount rate of 7% per year and an estimated life of 30 years were used.

Having much difficulty with this question, would appreciate some guidance with formulas, and explanation as to what is happening.
Answer #1

B/C ratio = 2, Initial Cost = $ 1.9 million

PV of Benefits = Initial Cost x (B/C) = 1.9 x 2 = $ 3.8 million

Discount Rate = 7 % and Tenure = 30 years

Let the benefits be $ m

Therefore, 3.8 = m x (1/0.07) x [1-{1/(1.07)^(30)}]

3.8 = m x 12.409

m = 3.8 / 12.409 = $ 306228.33

Benefit - Annual M&O Costs = 115000

Annual M&O Costs = 306228.33 - 115000 = $ 191228.33